# CRAFT Your Dataset: Task-Specific Synthetic Dataset Generation Through Corpus Retrieval and Augmentation

### v1

@@ Abstract
Building high-quality datasets for specialized tasks is a time-consuming and resource-intensive process that often requires specialized domain knowledge. We propose Corpus Retrieval and Augmentation for Fine-Tuning (CRAFT), a method for generating synthetic datasets, given a small number of user-written few-shots that demonstrate the task to be performed. Given the few-shot examples, we use large-scale public web-crawled corpora and similarity-based document retrieval to find other relevant human-written documents. Lastly, instruction-tuned large language models (LLMs) augment the retrieved documents into custom-formatted task samples, which then can be used for fine-tuning. We demonstrate that CRAFT can efficiently generate large-scale task-specific training datasets for four diverse tasks: biology question-answering (QA), medicine QA and commonsense QA as well as summarization. Our experiments show that CRAFT-based models outperform or achieve comparable performance to general LLMs for QA tasks, while CRAFT-based summarization models outperform models trained on human-curated data by 46 preference points.
@@ 1 Introduction
Traditionally, creating high-quality datasets for specific tasks involves a time-consuming and resource-intensive process, often requiring extensive manual curation and annotation (e.g. Marcus et al. (1993)). This challenge is particularly acute for low-resource domains or novel tasks where existing datasets may be limited or non-existent.
@@
We demonstrate the effectiveness of CRAFT on four diverse tasks: three QA tasks – in biology, medicine and commonsense – as well as a text summarization generative task. Our results show that models fine-tuned on CRAFT-generated datasets achieve performance that is either better than or comparable to instruction-tuned LLMs. This holds across diverse tasks, LLMs, and dataset sizes, highlighting the effectiveness of our approach. We publicly release the code to craft datasets for other tasks as well as all datasets and checkpoints at github.com/ziegler-ingo/CRAFT.
@@ 2.1 Optimizing LLMs for Specific Tasks
Instruction tuning (Wei et al., 2022) is a type of full fine-tuning that optimizes a model to produce more relevant answers to questions or instructions (Leike et al., 2018; Askell et al., 2021). This approach enables language models to better understand and follow user intents rather than simply continuing the input text. Instruction-tuned models regularly produce answers that are preferred by humans for tasks ranging from question-answering to summarization (Ouyang et al., 2022). The main challenge is to obtain a large high-quality dataset that is both task-specific and in the desired instruction-output format.
@@ 2.2 Synthetic Data Generation
A dataset is fully synthetic if the question or instruction, the possible context, as well as the answers are generated synthetically. For instance, Self-Instruct (Wang et al., 2023b), Unnatural Instructions (Honovich et al., 2023), Alpaca (Taori et al., 2023), and Evol-Instruct (Xu et al., 2023) are examples of fully synthetic general-purpose data generated by LLMs. More focused approaches for task-specific fine-tuning data generation have also been proposed, especially based around the rephrasing of already existing task-specific datasets (Yin et al., 2023; Gandhi et al., 2024). Methods that use general-purpose corpora have recently been proposed for generating pre-training data (Maini et al., 2024). When used for fine-tuning data generation, these methods are either based around complex and resource-intensive multi-agent workflows (Mitra et al., 2024) or are restricted to a small set of tasks, as the generation process relies on a model that has been fine-tuned for those tasks (Nayak et al., 2024).
The two greatest drawbacks of current approaches to fully synthetic data generation are repetition and low quality. Unnatural Instructions reported that a majority of their samples have a BERTScore (Zhang et al., 2020) of above 45% when compared to other samples in the generated dataset. Self-Instruct faces similar issues, with generated instructions often having ROUGE-L scores (Lin, 2004) greater than 0.4 compared to the provided seed instructions. Both approaches also only contain about 54%-56.5% correct samples, while the correctness rate in Alpaca is as low as 17% (Chen et al., 2024a). This suggests that a large portion of the samples in these datasets may not be useful for fine-tuning models.
@@
One recent approach is reverse instruction generation (Köksal et al., 2023), where a language model, provided with a human-curated output in context, generates the instruction that would have prompted this output. This produces more coherent and correct input-output pairs because the LLM does not need to generate the longer and more complex component of the data sample. There are also approaches where, conversely, the output is synthetically generated from human-curated input samples. Such methods employ distillation techniques to extract patterns from larger, more capable models to teach those patterns and skills to smaller models (Mukherjee et al., 2023; Mitra et al., 2023).
Partially synthetic data generation alleviates some of the quality and diversity concerns of fully synthetic data generation. However, taking a raw corpus document as the output can result in generating noisy or unnecessary information (Agarwal et al., 2007), but data augmentation can mitigate these problems when generating pre-training data (Maini et al., 2024). However, when data augmentation was used to generate fine-tuning data, it required GPT-4 (OpenAI, 2023) to build an intermediate synthetic dataset to fine-tune a sample creator model (Chen et al., 2024b). This approach can result in a sample creator model that is a distilled version of the larger model’s knowledge and data, while also limiting the model’s task flexibility, depending on the synthesized training data.
In contrast, CRAFT produces fully synthetic data but leverages the quality and diversity advantages of human-written documents from partially synthetic data generation approaches while removing noise through augmentation. Our approach does not require intermediate datasets, nor a separately fine-tuned model, nor knowledge distillation from a larger model; instead, it relies only on a small number of human-curated examples, retrieval, and in-context learning.
@@ 3.2 Few-Shot Examples
A small number of human-curated few-shots serve as the “definition” of the task, i.e., they indicate how the task is to be performed. The few-shot samples consist of three elements: (i) a long text that mirrors in language, content, and accuracy what a high-quality corpus sample from the web should look like, (ii) a natural language instruction for the task to be performed, which can take the form of a direct instruction or a question about the text, and (iii) an output that satisfies the instruction or answers the question the way that the final model should later respond. Length statistics for texts, instructions, and outputs of our few-shots can be found in the XS row of Appendix C.
@@ 3.4 Document Retrieval
Our retrieval system is task agnostic, both in terms of domain and complexity, in contrast to previous approaches (Ein-Dor et al., 2020; Dai et al., 2022; Lewis et al., 2021). The CRAFT approach relies on human-curated few-shot examples as query documents and can dynamically retrieve any document of the base corpora. As the few-shot samples include a text containing the domain, the instruction or question, as well as the output, the resulting embedding representation of the sample contains contextualized (Reimers and Gurevych, 2019) semantic information about both the domain and the nature of task to be performed. Relevant text documents that contain similar latent features as the few-shots are retrieved from the corpora by calculating similarity scores based on the embedded few-shots and corpus samples.
@@ 3.5 Task Sample Synthesis
The retrieved documents naturally contain noise (Agarwal et al., 2007) and lack the formatting required for fine-tuning. Therefore, it is necessary to convert these free-text documents into appropriate task samples by removing noise and undesired sections.
To address this, we utilize instruction-tuning prompt templates (Sanh et al., 2022; Maini et al., 2024) to augment the free-text documents into task-specific training data while simultaneously eliminating noise. A few-shot task template consists of three elements: (i) one or more few-shots, (ii) a corpus sample, (iii) and a brief instruction for the model to generate instruction-output pairs from the content of the corpus sample. Aside from the brief instruction, it is easy to assemble these templates from material we already have. The template only structures all information from the instruction, the few-shots, and the retrieved corpus samples to generate one continuous string that serves as input for the model generating the synthetic task samples. In this setup, the contents of the few-shots serve as in-context examples for the completion of the instruction. Figure 2, step 3, shows an example of how these templates guide the model in augmenting the corpus samples into synthetic task samples. Any instruction-tuned language model can be used for this purpose. This augmentation step not only rephrases the text but also condenses the retrieved document down to the essential information required for the task. The result of this step produces the final synthetic instruction-output pairs that can be used to fine-tune a language model. Figure 2, step 4, shows an actual example output from the generated pool of synthetic training samples, and Appendix C provides an overview of length statistics from the stages of corpus retrieval up to the synthesized input-output pairs.
@@ 4 Experimental Setup
4.1 Tasks
We generate datasets for all tasks in sizes 100, 500, 5000, and 25,000. We refer to the few-shots as a dataset of size XS, and to the sizes ranging from 100 to 25,000 as S, M, L, and XL, respectively. Implementation details related to fine-tuning can be found in Appendix A.6.
Multiple-Choice QA:
@@ 4.1 Tasks
We generate three synthetic QA datasets for biology (BioQA), medicine (MedQA), and commonsense (CSQA). The datasets all follow the MMLU multiple-choice format (Hendrycks et al., 2021), where each question is accompanied by a number of answer options. Exactly one of these options is correct, and the task is to identify the correct answer. The expected output is the corresponding letter label of the correct answer.
Generative:
@@
In addition, we develop two synthetic datasets for the generative tasks of summarization and recipe generation (RecipeGen). The goal of summarization is to convey accurate and concise information; recipe generation is additionally focused on creating coherent and structured text that adheres to specific formatting and stylistic conventions (Wang et al., 2022). To build a synthetic summarization dataset, we first select a corpus sample and instruct the model to extract an extensive but unsummarized section of text. In the second step, the extracted section is transformed into a summary format, optionally incorporating elements from the raw text, such as abstracts, conclusions, or TLDRs. This approach avoids using the original corpus samples as to-be-summarized text documents, which can be lengthy and overly broad and may result in uninformative summaries.

### v2

@@ Abstract
Building high-quality datasets for specialized tasks is a time-consuming and resource-intensive process that often requires specialized domain knowledge. We propose Corpus Retrieval and Augmentation for Fine-Tuning (CRAFT), a method for generating synthetic datasets, given a small number of user-written few-shots that demonstrate the task to be performed. Given these examples, CRAFT uses large-scale public web-crawled corpora and similarity-based document retrieval to find other relevant human-written documents. Lastly, instruction-tuned large language models (LLMs) augment the retrieved documents into custom-formatted task samples, which then can be used for fine-tuning. We demonstrate that CRAFT can efficiently generate large-scale task-specific training datasets for four diverse tasks: biology, medicine, and commonsense question-answering (QA), as well as summarization. Our experiments show that CRAFT-based models outperform or match general LLMs on QA tasks, while exceeding models trained on human-curated summarization data by 46 preference points. CRAFT outperforms other synthetic dataset generation methods such as Self- and Evol-Instruct, and remains robust even when the quality of the initial few-shots varies.
@@ 1 Introduction
Traditionally, creating high-quality datasets for specific tasks involves a time-consuming and resource-intensive process, often requiring extensive manual curation and annotation (e.g., Marcus et al. (1993)). This challenge is particularly acute for low-resource domains or novel tasks where existing datasets may be limited or non-existent.
@@
We demonstrate the effectiveness of CRAFT on four diverse tasks: three QA tasks – in biology, medicine and commonsense – as well as a text summarization generative task. Our results show that models fine-tuned on CRAFT-generated datasets achieve performance that is either better than or comparable to instruction-tuned LLMs. Moreover, CRAFT not only outperforms other fully synthetic data generation methods, such as Self-Instruct (Wang et al., 2023) and Evol-Instruct (Xu et al., 2023), but also exhibits robustness to variations in the quality of the initial few shots. This holds across diverse tasks, LLMs, and dataset sizes, highlighting the effectiveness of our approach. We publicly release the code to craft datasets for other tasks as well as all datasets and checkpoints at github.com/ziegler-ingo/CRAFT.
@@ 2.1 Optimizing LLMs for Specific Tasks
Instruction tuning (Wei et al., 2022) is a type of full fine-tuning that optimizes a model to produce more relevant answers to questions or instructions (Leike et al., 2018; Askell et al., 2021). This approach enables language models to understand and follow user intents rather than simply continuing the input text. Instruction-tuned models produce answers preferred by humans for tasks ranging from question-answering to summarization (Ouyang et al., 2022). The challenge is obtaining a large high-quality dataset that is both task-specific and in the desired instruction-output format.
@@ 2.2 Synthetic Data Generation
A dataset is fully synthetic if the question or instruction, the possible context, as well as the answers are generated synthetically. Methods such as Self-Instruct (Wang et al., 2023) and Evol-Instruct (Xu et al., 2023) generate general-purpose datasets by prompting LLMs to create task instructions and corresponding outputs. Other approaches focus on generating task-specific fine-tuning data through the rephrasing of existing datasets (Yin et al., 2023; Gandhi et al., 2024) or on synthesizing pre-training data from general-purpose corpora (Maini et al., 2024). When applied to fine-tuning, these methods are either implemented via complex, resource-intensive multi-agent workflows (Mitra et al., 2024) or are constrained to a narrow set of tasks because the generation process relies on models fine-tuned for those specific tasks (Nayak et al., 2024).
Two primary challenges of current approaches to fully synthetic data generation are repetition and inconsistent data quality. Evaluations have indicated that many samples in fully synthetic datasets tend to exhibit high similarity to one another or to the seed samples (Honovich et al., 2023; Wang et al., 2023). Moreover, reported correctness rates for these datasets suggest that a substantial portion of the generated samples may not meet desired correctness standards (Chen et al., 2024a). These observations highlight areas where fully synthetic data generation can be refined.
@@
One recent approach is reverse instruction generation (Köksal et al., 2023), where a language model, provided with a human-curated output in context, generates the instruction that would have prompted this output. This produces more coherent and correct input-output pairs because the LLM does not need to generate the longer and more complex component of the data sample. There are also approaches where, conversely, the output is synthetically generated from human-curated input samples. Such methods employ distillation to extract patterns from larger models to teach those patterns and skills to smaller models (Mukherjee et al., 2023; Mitra et al., 2023).
Partially synthetic data generation can alleviate some of the quality and diversity concerns inherent to fully synthetic approaches. However, using a raw corpus document as the output may introduce noisy or unnecessary information (Agarwal et al., 2007). Data augmentation has been shown to mitigate these issues for pre-training data generation (Maini et al., 2024). When applied to fine-tuning data, though, such augmentation typically requires a powerful model (e.g., GPT-4 (OpenAI, 2023)) to construct an intermediate synthetic dataset for fine-tuning a sample creator model (Chen et al., 2024b). This multi-stage approach can lead to a sample creator model that essentially distills the larger model’s knowledge, potentially limiting task flexibility based on the synthesized training data. Alternatively, some methods have started to use retrieval to enhance diversity: SynthesizRR (Divekar and Durrett, 2024) uses knowledge distillation from larger models while incorporating retrieved documents as in-context examples. RADA (Seo et al., 2024) relies on already structured input-output pairs from other existing datasets and retrieves them to improve the generation prompt for low-resource tasks.
In contrast, CRAFT streamlines this process by operating directly on unstructured, free-text corpora. Our approach produces fully synthetic data but leverages the quality and diversity advantages of human-written documents from partially synthetic data generation approaches while removing noise through augmentation. CRAFT does not require intermediate datasets, nor a separately fine-tuned model, nor knowledge distillation from a larger model; instead, it relies only on a small number of human-curated examples, retrieval, and in-context learning.
@@ 3.2 Few-Shot Examples
A small number of human-curated few-shots serve as the “definition” of the task, i.e., they indicate how the task is to be performed. The few-shot samples consist of three elements: (i) a long text that mirrors in language, content, and accuracy what a high-quality corpus sample from the web should look like, (ii) a natural language instruction for the task to be performed, which can take the form of a direct instruction or a question about the text, and (iii) an output that satisfies the instruction or answers the question the way that the final model should later respond. Length statistics for texts, instructions, and outputs of our few-shots can be found in the XS row of Appendix E.
@@ 3.4 Document Retrieval
Our retrieval system is task agnostic, both in terms of domain and complexity, in contrast to previous approaches (Ein-Dor et al., 2020; Dai et al., 2022; Lewis et al., 2021). The CRAFT approach relies on human-curated few-shot examples as query documents and can dynamically retrieve any document of the base corpora. As the few-shot samples include a text containing the domain, the instruction or question, as well as the output, the resulting embedding representation of the sample contains contextualized (Reimers and Gurevych, 2019) semantic information about both the domain and the nature of the task to be performed. Relevant text documents that contain similar latent features as the few-shots are retrieved from the corpora by calculating similarity scores based on the embedded few-shots and corpus samples.
@@ 3.5 Task Sample Synthesis
The retrieved documents naturally contain noise (Agarwal et al., 2007) and lack the formatting required for fine-tuning. Therefore, it is necessary to convert these free-text documents into appropriate task samples by removing noise and undesired sections. To address this, we utilize instruction-tuning prompt templates (Sanh et al., 2022; Maini et al., 2024) to augment free-text documents into task-specific training data while eliminating noise. A few-shot task template consists of three elements: (i) one or more few-shots, (ii) a corpus sample, and (iii) a brief instruction for the model to generate instruction-output pairs from the corpus sample. The template structures all information from the instruction, the few-shot examples, and the retrieved corpus sample, which together serves as input for the model that generates synthetic task samples.
Figure 2, step 3, shows an example of how these templates guide the model in augmenting the corpus samples into synthetic task samples. This augmentation step not only rephrases the text but also condenses the retrieved document down to the essential information required for the task. This step produces final synthetic instruction-output pairs that can be used to fine-tune a language model. Figure 2, step 4, shows an actual example output from the generated pool of synthetic training samples, and Appendix E provides an overview of length statistics from the stages of corpus retrieval up to the synthesized input-output pairs.
@@ 4 Experimental Setup
This section summarizes how we implemented the CRAFT pipeline, the tasks on which we evaluate, dataset details for both CRAFT and the human-annotated baselines, model baselines and training details, as well as the evaluation metrics.
4.1 CRAFT Implementation Overview
We construct the embedding database from four large-scale corpora: C4 (Raffel et al., 2020), English Wikipedia, Stack Exchange (Flax Sentence Embeddings Team, 2021), and WikiHow (Koupaee and Wang, 2018). After filtering documents by length (200-25,000 characters), this collection comprises 383 million documents. We generate 384D embeddings for each document using SentenceTransformers (Reimers and Gurevych, 2019) with MiniLM (Wang et al., 2020) version multi-qa-MiniLM-L6-cos-v1.
During the retrieval phase, cosine similarity is computed between each few-shot example and the documents in each corpus shard (approx. 350K documents), retaining the top 5% from each shard. Then, a global top-k selection was performed over the retained candidates. We employ a mixed strategy to balance topic coverage and specificity: 50% of documents are retrieved based on their top-k cosine similarity to individual few-shot examples, while the other 50% are retrieved based on similarity to the averaged embedding of all few-shots.
For task samples synthesis, we use Mistral 7B Instruct v0.2 (Jiang et al., 2023) with temperature=0.7, top-k=40 (Fan et al., 2018), and top-p=0.9 (Holtzman et al., 2020). Input prompts used three randomly sampled few-shots. To ensure data quality and structural integrity, we generate outputs as JSON objects, which allows for format validation. Finally, we enhance dataset diversity by performing a deduplication step, filtering out generated samples that are highly similar to the seed few-shots or to each other using fuzzy string matching (Ranjith et al., 2022) with a token set similarity ratio of 0.85. Further implementation details and filtering statistics are provided in Appendices A and D, respectively.
4.2 Tasks and Datasets
To test the performance of the CRAFT pipeline, we evaluate it on five different tasks: multiple-choice (MC) biology QA, MC commonsense QA, MC medicine QA, summarization, and recipe generation. We first describe the details of CRAFT and then introduce the datasets for the human-annotated baselines.
4.2.1 CRAFT Datasets
We generate datasets for all tasks with sizes of 100, 500, 5,000, and 25,000. We refer to the few-shot datasets (with 8 or 32 examples) as size XS, and to the datasets with 100, 500, 5,000, and 25,000 examples as sizes S, M, L, and XL, respectively. For few-shot curation, we do not refer to any existing datasets; instead, the authors manually curate examples from corpora (see Appendix A.1 for sources used).
@@ 4.2.1 CRAFT Datasets
Multiple-Choice QA: We generate three synthetic QA datasets in the domains of biology (BioQA), commonsense (CSQA), and medicine (MedQA). All datasets follow the MMLU multiple-choice format (Hendrycks et al., 2021), where each question is accompanied by several answer options. Exactly one of these options is correct. The task is to output the correct answer, the letter label corresponding to the correct option.
Generation: We develop two synthetic datasets for the generation tasks of recipe generation (RecipeGen) and summarization. The goal of summarization is to convey accurate and concise information, while recipe generation focuses on creating coherent and structured text that adheres to specific formatting and stylistic conventions (Wang et al., 2022).
@@
To build a synthetic summarization dataset, we first select a corpus sample and instruct the model to extract an extended section of text. In the second step, the extracted section is transformed into a summary, optionally incorporating elements from the raw text (e.g., ‘abstract:’, ‘conclusion:’, or ‘TLDR’). This approach avoids using full corpus samples as the text to be summarized, as they can be lengthy and overly broad, potentially resulting in uninformative summaries. For recipe generation, our goal is to generate a list of ingredients and cooking steps for a specific recipe.
4.2.2 Human-Annotated Datasets
We select human-annotated datasets as optimal baselines for each task to compare the performance of models trained on CRAFT. For evaluation, we use the test split of these datasets, as detailed in §4.5; therefore, the human-annotated datasets can also be considered in-domain.
BioQA: We use the biology subset of ScienceQA (Lu et al., 2022) with 1,192 training samples without images. ScienceQA sourced expert-curated question-answer pairs from online learning platforms, ensuring high quality and accuracy. The dataset’s answer options range from two to five, include a single correct answer per question, and are randomized to prevent pattern recognition.
MedQA: We use MedMCQA (Pal et al., 2022) and randomly select 25,000 samples from the training split. The dataset comprises entrance exam questions from two of India’s postgraduate institutions. All samples come from preparation materials or real exams created by medical professionals, with each question containing four answer options and one correct answer.
CSQA: We use CommonsenseQA 2.0 (Talmor et al., 2021) and select 9,264 samples from the training split. The dataset was generated through a gamified yet controlled question generation process where players earned points by designing challenging yes/no questions that outperform an AI model. Generated questions were validated by other players, independent validators, and another model to ensure they were well-formed, answerable, and representative of common sense.
RecipeGen: We use RecipeNLG (Bień et al., 2020) and select 25,000 samples from the training split. The recipes were scraped from cooking websites and post-processed through fine-grained cleaning and formatting to ensure correctness. Each recipe includes a title, ingredient list, and cooking steps. We exclude samples present in C4 based on provided URLs.
Summarization: We use CNN-DailyMail (See et al., 2017) and randomly select 25,000 samples from the training split. This dataset is commonly used for summarization tasks as its CNN and DailyMail articles have highlights presented in abstract or bullet point formats.
4.3 Baselines
We compare CRAFT models trained on synthetic data against multiple baselines:
Few-shot is fine-tuned only on XS-size CRAFT datasets containing human-curated few-shot examples. This serves as our primary baseline, representing all human-curated data in our pipeline.
Instruct: Mistral 7B Instruct v0.2 (Jiang et al., 2023) is instruction-tuned on proprietary instruction-following datasets. This provides a meaningful comparison to a similarly-sized instruction-tuned model, though trained on undisclosed data of unknown quality and quantity. Surpassing this baseline would indicate CRAFT’s ability to generate high-quality synthetic data.
In-Domain (ID) is fine-tuned on human-curated training splits from evaluation datasets. This baseline represents the optimal performance achievable with human quality datasets.
We additionally compare CRAFT against two popular synthetic data generation methods:
Self-Instruct (Wang et al., 2023) distills new task samples from instruction-tuned models, typically using fewer than 10 seed examples per task.
Evol-Instruct (Xu et al., 2023) generates new samples by rewriting seed examples to be more complex, to incorporate more domain-specific concepts, or to add step-by-step reasoning.
We initialize these methods with 8 few-shot examples to follow Self-Instruct’s setup and also generate four dataset sizes (S: 100, M: 500, L: 5,000, XL: 25,000) to follow CRAFT’s approach. We use each method’s original hyperparameters and the same seed examples.
4.4 Training and Optimization
We fine-tune base models using either CRAFT datasets or human-annotated datasets. Models fine-tuned with human-annotated datasets are denoted with the suffix -ID (in-domain), while models fine-tuned with CRAFT datasets are labeled as -CRAFT#, where # represents the training split size (XS, S, M, L, or XL). Our primary experimental setting uses Mistral 7B v0.2 as the base model with 32 few-shot examples in CRAFT. Additional experiments may use 8-shot settings or another base model, Llama 3 8B, specified explicitly when used. We also evaluate (as baselines) instruct versions of base models without fine-tuning.
For all experiments, low-rank adaptation (Hu et al., 2021, LoRA) fine-tuning is performed using 16-bit BrainFloat (Abadi et al., 2016) as the computation type. All implementations use PyTorch (Paszke et al., 2019) and HuggingFace libraries (Wolf et al., 2020). For optimization, the adaptive momentum optimizer with decoupled weight decay (Loshchilov and Hutter, 2019) of 5% and a learning rate of $1\times 10^{-4}$ is employed. A linear learning rate schedule is applied with a warm-up ratio of 10%. Models are fine-tuned for three epochs across all tasks and dataset sizes for CRAFT. When training only on human-curated few-shots, we adopt a batch size of 2; otherwise, we use batch size 16 or equivalent gradient accumulation steps. Following Dettmers et al. (2024), we apply LoRA adapters to every linear layer (query-, key-, value-, output-, gate-, up- and down-projection matrices) with rank 64 and $\alpha$ 64 and 0.1 dropout rate. Bias terms in update matrices are deactivated. This configuration adds 2.3% (160 million parameters) to the base model’s 7 billion parameters as LoRA adapters. Frozen base parameters remain unchanged during training, with updated parameters merged post-training.
Figure 3: Performance scaling with increasing data size across multiple tasks using CRAFT with 32 few-shot examples. Graphs demonstrate consistent improvements as training data grows from few-shot (XS) to 25,000 synthetic samples (XL). CRAFT models consistently match or exceed Instruct performance (dotted red line). Shaded regions indicate standard deviation across three runs.